# Sensitive Content Classification in Social Media: A Holistic Resource and Evaluation

### v2

@@ Abstract
The detection of sensitive content in large datasets is crucial for ensuring that shared and analysed data is free from harmful material. However, current moderation tools, such as external APIs, suffer from limitations in customisation, accuracy across diverse sensitive categories, and privacy concerns. Additionally, existing datasets and open-source models focus predominantly on toxic language, leaving gaps in detecting other sensitive categories such as substance abuse or self-harm. In this paper, we put forward a unified dataset tailored for social media content moderation across six sensitive categories: conflictual language, profanity, sexually explicit material, drug-related content, self-harm, and spam. By collecting and annotating data with consistent retrieval strategies and guidelines, we address the shortcomings of previous focalised research. Our analysis demonstrates that fine-tuning large language models (LLMs) on this novel dataset yields significant improvements in detection performance compared to open off-the-shelf models such as LLaMA, and even proprietary OpenAI models, which underperform by 10-15% overall. This limitation is even more pronounced on popular moderation APIs, which cannot be easily tailored to specific sensitive content categories, among others. Equal contribution.
@@ A Holistic Resource and Evaluation
Dimosthenis Antypas*1, Indira Sen*2, Carla Perez-Almendros1 Jose Camacho-Collados1, Francesco Barbieri 1 Cardiff NLP, Cardiff University, United Kingdom 2 University of Mannheim 1 {AntypasD,CamachoColladosJ,perezalmendrosc}@cardiff.ac.uk 2 indira.sen@uni-mannheim.de
@@ 1 Introduction
Consider the case of a researcher or a data analyst who needs to filter sensitive content from a large dataset. Such a task is crucial to ensure that data shared or analysed does not include harmful or inappropriate material. One might initially consider using external tools like Perspective [^1] https://perspectiveapi.com/ or OpenAI moderation APIs [^2] https://platform.openai.com/docs/guides/moderation to assess and filter sensitive content. However, this approach often falls short, presenting important limitations for an effective identification of inappropriate content online Udupa et al. (2023). For instance, they usually offer limited customisation capabilities (e.g., how can the model be improved if it fails on specific domains or keywords?), and limited sensitive categories coverage (lacking in detecting self-harm Uban and Rosso (2020) for example). Finally, these tools rely on external servers, which raises concerns about data privacy and security Oseni et al. (2021); Gupta et al. (2023).
@@
Alternatively, one might consider using existing datasets and open source models for sensitive content detection. This could be a viable option if the primary focus was on detecting toxic language, given the abundance of resources available in this area. However, if the goal extends to identifying additional sensitive categories such as sexually explicit content, drugs, self-harm, and spam, the situation becomes more challenging. Data on these less-explored categories is limited and sometime outdated. For instance, those categories could be covered on dataset that are biased Wiegand et al. (2019), old or inaccessible even in an anonymized manner Tadesse et al. (2019); Sawhney et al. (2018), too small-scaled or rely on an handful of keywords to extract the data Ding et al. (2016). This limited approach can result in incomplete or less accurate detection of sensitive content.
@@
Existing solutions either require sending data to external servers or fail to address the full spectrum of sensitive content categories. In response to these challenges, this paper proposes a new holistic approach: a unified dataset for detecting sensitive text across a broad range of categories, including (1) conflictual language, (2) profanity, (3) sexually explicit material, (4) drug-related content, (5) self-harm, and (6) spam [^3] We use social media platform user guidelines to be sure to have reliable guidelines.. This dataset can then be used for both evaluation and fine-tuned model to all these categories under a single framework.
@@
Our approach involves collecting and re-annotating data to ensure consistent quality across sensitive classes. The alternative of putting together a collection of existing datasets to create one single dataset would include several limitations as (1) there would be a bias towards the data distribution (different retrieval strategies, topics, source platforms), (2) annotation guidelines and quality would differ, (3) each text would include only one sensitive dimension (even if the text includes multiple sensitive categories).
@@
  • We introduce the X-Sensitive dataset, manually annotated and tailored for social media content, featuring multiple categories and designed to be resilient against keyword and domain shifts.
  • Study the interplay between sensitive categories and how these categories vary across different annotator demographics.
  • Analyse the performance of current models: The best results are achieved from large language models (LLMs) with 8 billion parameters, fine-tuned on our dataset. However, smaller language models [^4] Pre-trained on social media language. (355 million parameters) show only about 2% less accuracy.
@@ 2 Related Work
Our current work aims to bridge the gap between current academic research in content moderation and the needs of content moderators in realistic scenarios. While hate speech and toxic language are widely studied in NLP, there is little research on other types of sensitive content that platforms seek to detect and moderate, such as sexually explicit content or content about illicit substances Arora et al. (2023). To that end, our work in situated at the intersection of NLP research on harmful language detection and research on platform governance and content moderation.
@@ 3.1 Taxonomy
Pornographic or other types of sexual content. We collect and download 50+ textual abusive language datasets from hatespeechdata.com. We then use the Perspective API to label these datasets with the ‘sexually explicit’ endpoint and then use the labeled data for finetuning the XLM-T sexually explicit content classifier.
@@
Attacks against protected attributes like race, color, caste, ethnicity, national origin, religion, sexual orientation, gender identity, disability, or veteran status, immigration status, socio-economic status, age, weight or pregnancy status.
@@ 3.2 Tweet Collection
For self-harm and spam, unigrams are not as informative as they are for other categories. Therefore for the former, we look at past research on eating disorders and suicidal ideation Chancellor et al. (2016b, 2021) and obtain phrases (ngrams with a high TF-IDF score) from the Reddit data used to train the self-harm phase 1 classifier. For spam, we use the dataset from Founta et al. (2018), particularly the tweets that were labelled containing spam and obtain ngrams from there using a similar method. We manually assess each of the keywords for all categories and remove low precision words like ‘snow’ for drugs. While snow may refer to cocaine in some contexts, most tweets containing it do not use it in that sense. After this manual inspection, we apply our cluster-based keyword expansion technique. We again manually assess the keywords and include only those that are relevant to the category. The final statistics of our keywords are listed in Appendix A, Table 4.
@@ 3.4 Statistics
Our dataset displays a skewed distribution of classes as seen if Table 1 with profanity being the most populated class present (30.4%). This uneven distribution represents a realistic representation of sensitive content in social media as seen in previous similar studies Beknazar-Yuzbashev et al. (2022) where is estimated a 5% - 7% of content displayed is inappropriate, making X-Sensitive ideal for usage in real world applications.
@@
In general differences between the classes are revealed even when looking at basic statistics such as the average length of tweets and the presence of emojis in them. As seen in Table 1 tweets labelled as spam tend to be longer on average and include a higher number of emojis, characteristics frequently found on spam messages Robinson and Mago (2022). Similarly, a higher usage of emojis is observed in tweets flagged as sexual explicit content, as specific emojis are often used as representation of sexual acts Thomson et al. (2018). Furthermore, when examining the top terms of each class based on lexical specificty scores Camacho-Collados et al. (2016), a clear distinction between the classes is observed, which also serves as a sanity check for the quality of our dataset.
@@ 4 Experimental Setting
In this section, we set out the common experimental framework which serve for the basis of the evaluation.
@@ 4.1 Data and Settings
To evaluate X-Sensitive and establish baselines of it’s difficulty we establish two distinct settings: binary and multi-label classification. In the binary setting, tweets will be classified into one of two categories, distinguishing between sensitive and not sensitive content. This approach simplifies the classification process, focusing on the presence or absence of sensitive characteristics in general. In the multi-label setting, tweets can belong to multiple sensitive categories simultaneously, allowing for a more finegrained analysis that captures the complexity of the content. This dual approach enables a comprehensive evaluation of our dataset’s versatility and the classifier’s robustness in handling varying degrees of complexity in sensitive content detection.
@@ 4.2 Comparison Systems
For the evaluation we are interested in comparing three types of approaches: finetuning on the same dataset (Section 4.2.1), LLMs with in-context learning either zero- or few-shot (Section 4.2.2), and out of the box content moderation systems (Section 4.2.3). All the systems are clearly not fully comparable, but our dataset can serve as the basis for establishing this basic ground comparison.
4.2.1 Finetuning
We evaluate three distinct models tailored for various applications, including general-purpose and those specialized for social media, each differing in size for our fine-tuning experiments. The large version of RoBERTa Liu (2019) is tested in order to assess the performance of smaller, non specialised, masked language models on our dataset. Time-lm, tlm, Loureiro et al. (2022), a RoBERta based model trained on a large X corpus of 154 million tweets is also evaluated to assess the performance of specialised models on social media. The two models are fine-tuned using the implementations provided by Hugging Face Wolf et al. (2020) and optimising hyper parameters (learning rate, training epochs, warmup steps) is conducted using Ray Tune Liaw et al. (2018) [^6] Details of the models used can be found in Appendix C.. Finally, the 8 billion version of Llama-3, Llama3-8b, AI@Meta (2024) is also fine-tuned on our dataset by utilising quantisation and PEFT Liu et al. (2021); Mangrulkar et al. (2022) explore the capabilities of more recent and larger-scale models.
@@ 4.2.2 Zero- and Few-shot
gpt-4o: the currently latest model from OpenAI which significantly outperforms it’s predecessor.
Assessing the performance in zero- and few-shot settings, help us to exlore the capabilities and limitations of these large language models for sensitive content detection.
Training Model Binary multi-label Conflictual Profaninty Sex Drugs Selfharm Spam Not sens.
@@ 4.2.3 Out of the box Systems
Meta-Llama-Guard-2-8B (llama-g) Inan et al. (2023) is a specialised version of LLama-3 that aims to classify content based on a safety risk taxonomy of 11 harm categories Vidgen et al. (2024)\@footnotemark. For our use case we consider only 5 of the categories that correspond better to the taxonomy used in X-Sensitive. Specifically we consider: "Hate" for Conflictual, "Suicide & Self-Harm" for self harm; and "Sexual Content", "Sex-Related Crimes", and "Child Sexual Exploitation" for sexual explicit content.
@@ 5.1 Fine-tuned Systems
All the fine-tuned models demonstrate high performance with RoBERTa as the least effective, achieving macro-F1 scores of 82.4 in the binary setting and 64.7 in the multi-label setting. The specialised training corpus of tlm appears to enhance its performance, as it consistently surpasses RoBERTa of the same architecture in both settings. Moreover, the fine-tuned version of the larger and more recent llama3-8b model achieves the best overall results in both settings, with macro-F1 scores of 85.6 in the binary setting and 69.8 in the multi-label setting, notably achieved without any hyper-parameter tuning, unlike the other models. Overall, the fine-tuned models tend to struggle the most with the least represented classes, such as Drugs, Selfharm, and Spam. Interestingly, despite comprising 17.3% of the total entries, the models under perform in the Conflictual category, while they exhibit better performance in the less prevalent Sexual Explicit Content class, which accounts for only 9.7% of tweets. This disparity may indicate the models’ difficulties in identifying subtler features within the Conflictual category.
@@ 8 Limitations
Finally, while we conduct an in-depth analysis using the results of six different models, there is significant room for improvement in terms of analysis and model development. This includes, investigating the performance of models of different architectures and optimising the prompts used.
@@ 9 Ethics Statement
We also ensure fair treatment of the annotators who labelled the dataset by: 1) compensating them fairly at an average rate of 12\$ per hour, and 2) not sharing or storing personal identification information. As annotator demographics ply an important role in the perception of toxicity, following Prabhakaran et al. (2021), we will release the data, disaggregated by individual annotator labels, while making sure that the demographic information is course enough to prevent deanonymization of the crowdworkers.
@@ Appendix A Data Collection
Table 4 displays the total number of keywords and see words used in our data collection.
@@ B.2 Annotator Agreement
Table 5: Krippendorff’s alpha withing each gender group of coders.
@@
Table 6: Krippendorff’s alpha withing each age group of coders.
@@
Age conflictual profanity sex drugs selfharm spam AVG Coders
@@
Gender conflictual profanity sex drugs selfharm spam AVG Coders
@@ C.2 Taxonomies
’self_harm’: ’Self Harm’,
’self_harm_instructions’: ’Self Harm’,
’self_harm_intent’: ’Self Harm’,
@@
’self-harm’: ’Self Harm’,
@@
’self-harm/intent’: ’Self Harm’,
’self-harm/instructions’: ’Self Harm’,
@@
C.3 Extended Results

### v3

@@ Abstract
The detection of sensitive content in large datasets is crucial for ensuring that shared and analysed data is free from harmful material. However, current moderation tools, such as external APIs, suffer from limitations in customisation, accuracy across diverse sensitive categories, and privacy concerns. Additionally, existing datasets and open-source models focus predominantly on toxic language, leaving gaps in detecting other sensitive categories such as substance abuse or self-harm. In this paper, we put forward a unified dataset tailored for social media content moderation across six sensitive categories: conflictual language, profanity, sexually explicit material, drug-related content, self-harm, and spam. By collecting and annotating data with consistent retrieval strategies and guidelines, we address the shortcomings of previous focalised research. Our analysis demonstrates that fine-tuning large language models (LLMs) on this novel dataset yields significant improvements in detection performance compared to open off-the-shelf models such as LLaMA, and even proprietary OpenAI models, which underperform by 10-15% overall. This limitation is even more pronounced on popular moderation APIs, which cannot be easily tailored to specific sensitive content categories, among others. Equal contribution. $\dagger$ Work done while at Snap Inc.
@@ A Holistic Resource and Evaluation
[…]
@@ 1 Introduction
Consider the case of a researcher or a data analyst who needs to filter sensitive content from a large dataset. Such task is crucial to ensure that data shared or analysed does not include harmful or inappropriate material. One might initially consider using external tools like Perspective [^1] https://perspectiveapi.com/ or OpenAI moderation APIs [^2] https://platform.openai.com/docs/guides/moderation to assess and filter sensitive content. However, this approach often falls short, presenting important limitations for an effective identification of inappropriate content online Udupa et al. (2023). For instance, they usually offer limited customisation capabilities (e.g., how can the model be improved if it fails on specific domains or keywords?), and limited sensitive categories coverage (lacking in detecting self-harm Uban and Rosso (2020), for example). Finally, these tools rely on external servers, which raises concerns about data privacy and security Oseni et al. (2021); Gupta et al. (2023).
@@
Alternatively, one might consider using existing datasets and open-source models for sensitive content detection. This could be a viable option if the primary focus was on detecting toxic language, given the abundance of resources available in this area. However, if the goal extends to identifying additional sensitive categories such as sexually explicit content, drugs, self-harm or spam, the situation becomes more challenging. Data on these less-explored categories is limited and sometimes outdated. For instance, those categories could be covered on datasets that are biased Wiegand et al. (2019), old or inaccessible even in an anonymized manner Tadesse et al. (2019); Sawhney et al. (2018), too small-scaled or rely on an handful of keywords to extract the data Ding et al. (2016). This limited approach can result in incomplete or less accurate detection of sensitive content.
@@
Existing solutions either require sending data to external servers or fail to address the full spectrum of sensitive content categories. In response to these challenges, this paper proposes a new holistic approach: a unified dataset for detecting sensitive text across a broad range of categories, including (1) conflictual language, (2) profanity, (3) sexually explicit material, (4) drug-related content, (5) self-harm, and (6) spam [^3] We use social media platform user guidelines to be sure to have reliable guidelines.. This dataset can then be used for both evaluation and fine-tuned models to all these categories under a single framework.
@@
Our approach involves collecting and re-annotating data to ensure consistent quality across sensitive classes. The alternative of putting together a collection of existing datasets to create one single dataset would include several limitations, as (1) there would be bias towards the data distribution (different retrieval strategies, topics, source platforms), (2) annotation guidelines and quality would differ, and (3) each text would include only one sensitive dimension (even if the text includes multiple sensitive categories).
@@
  • New dataset: We introduce the X-Sensitive dataset, manually annotated and tailored for social media content, featuring multiple categories and designed to be resilient against keyword and domain shifts.
  • Sensitive category analysis: We study the interplay between sensitive categories and how these categories vary across different annotator demographics.
  • Model evaluation: The best results are achieved from large language models (LLMs) with 8 billion parameters, fine-tuned on our dataset. However, smaller language models [^4] Pre-trained on social media language. (355 million parameters) show only about 2% less accuracy.
@@ 2 Related Work
Our current work aims to bridge the gap between current academic research in content moderation and the needs of content moderators in realistic scenarios. While hate speech and toxic language are widely studied in NLP, there is little research on other types of sensitive content that platforms seek to detect and moderate, such as sexually explicit content or content about illicit substances Arora et al. (2023). To that end, our work is situated at the intersection of NLP research on harmful language detection and research on platform governance and content moderation.
@@ 3.1 Taxonomy
Pornographic or other types of sexual content. We collect and download 50+ textual abusive language datasets from hatespeechdata.com. We then use the Perspective API to label these datasets with the ‘sexually explicit’ endpoint and then use the labelled data for fine-tuning the XLM-T sexually explicit content classifier.
@@
Attacks against protected attributes like race, colour, caste, ethnicity, national origin, religion, sexual orientation, gender identity, disability, or veteran status, immigration status, socio-economic status, age, weight or pregnancy status.
@@ 3.2 Message Collection
For self-harm and spam, uni-grams are not as informative as they are for other categories. Therefore for the former, we look at past research on eating disorders and suicidal ideation Chancellor et al. (2016b, 2021) and obtain phrases (n-grams with a high TF-IDF score) from the Reddit data used to train the self-harm phase 1 classifier. For spam, we use the dataset from Founta et al. (2018), particularly the tweets that were labelled containing spam and obtain n-grams from there using a similar method. We manually assess each of the keywords for all categories and remove low precision words like ‘snow’ for drugs. While snow may refer to cocaine in some contexts, most tweets containing it do not use it in that sense. After this manual inspection, we apply our cluster-based keyword expansion technique. We again manually assess the keywords and include only those that are relevant to the category. The final statistics of our keywords are listed in Appendix A, Table 4.
@@ 3.4 Statistics
Our dataset displays a skewed distribution of classes as seen if Table 1 with profanity being the most populated class present (30.4%). This uneven distribution represents a realistic representation of sensitive content in social media as seen in previous similar studies Beknazar-Yuzbashev et al. (2022) where it estimated a 5% - 7% of content displayed is inappropriate, making X-Sensitive ideal for usage in real world applications.
@@
In general differences between the classes are revealed even when looking at basic statistics such as the average length of tweets and the presence of emojis in them. As seen in Table 1 tweets labelled as spam tend to be longer on average and include a higher number of emojis, characteristics frequently found on spam messages Robinson and Mago (2022). Similarly, a higher usage of emojis is observed in tweets flagged as sexual explicit content, as specific emojis are often used as representation of sexual acts Thomson et al. (2018). Furthermore, when examining the top terms of each class based on lexical specificity scores Camacho-Collados et al. (2016), a clear distinction between the classes is observed, which also serves as a sanity check for the quality of our dataset.
@@ 4 Experimental Setting
In this section, we set out the common experimental framework which serve as the basis of the evaluation.
@@ 4.1 Data and Settings
To evaluate X-Sensitive and establish baselines of its difficulty we establish two distinct settings: binary and multi-label classification. In the binary setting, tweets will be classified into one of two categories, distinguishing between sensitive and not sensitive content. This approach simplifies the classification process, focusing on the presence or absence of sensitive characteristics in general. In the multi-label setting, tweets can belong to multiple sensitive categories simultaneously, allowing for a more fine-grained analysis that captures the complexity of the content. This dual approach enables a comprehensive evaluation of our dataset’s versatility and the classifier’s robustness in handling varying degrees of complexity in sensitive content detection.
@@ 4.2 Comparison Systems
For the evaluation, we are interested in comparing three types of approaches: fine-tuning on the same dataset (Section 4.2.1), LLMs with in-context learning either zero- or few-shot (Section 4.2.2), and out-of-the-box content moderation systems (Section 4.2.3). All the systems are clearly not fully comparable, but our dataset can serve as the basis for establishing this basic ground comparison.
4.2.1 Fine-tuning
We evaluate three distinct models tailored for various applications, including general-purpose and those specialized for social media, each differing in size for our fine-tuning experiments. The large version of RoBERTa Liu (2019) is tested in order to assess the performance of smaller, non specialised, masked language models on our dataset. Time-lm, tlm, Loureiro et al. (2022), a RoBERta based model trained on a large X corpus of 154 million tweets is also evaluated to assess the performance of specialised models on social media. The two models are fine-tuned using the implementations provided by Hugging Face Wolf et al. (2020) and optimising hyper parameters (learning rate, training epochs, warm-up steps) is conducted using Ray Tune Liaw et al. (2018) [^6] Details of the models used can be found in Appendix C.. Finally, the 8 billion version of Llama-3, Llama3-8b, AI@Meta (2024) is also fine-tuned on our dataset by utilising quantisation and PEFT Liu et al. (2021); Mangrulkar et al. (2022) explore the capabilities of more recent and larger-scale models.
@@ 4.2.2 Zero- and Few-shot
gpt-4o: the currently latest model from OpenAI which significantly outperforms its predecessor.
Assessing the performance in zero- and few-shot settings, helps us to exlore the capabilities and limitations of these large language models for sensitive content detection.
Training Model Binary multi-label Conflictual Profaninty Sex Drugs Self-harm Spam Not sens.
@@ 4.2.3 Out of the box Systems
Meta-Llama-Guard-2-8B (llama-g) Inan et al. (2023) is a specialised version of LLama-3 that aims to classify content based on a safety risk taxonomy of 11 harm categories Vidgen et al. (2024)\@footnotemark. For our use case we consider only 5 of the categories that correspond better to the taxonomy used in X-Sensitive. Specifically we consider: "Hate" for Conflictual, "Suicide & Self-Harm" for self-harm; and "Sexual Content", "Sex-Related Crimes", and "Child Sexual Exploitation" for sexual explicit content.
@@ 5.1 Fine-tuned Systems
All the fine-tuned models demonstrate high performance with RoBERTa as the least effective, achieving macro-F1 scores of 82.4 in the binary setting and 64.7 in the multi-label setting. The specialised training corpus of tlm appears to enhance its performance, as it consistently surpasses RoBERTa of the same architecture in both settings. Moreover, the fine-tuned version of the larger and more recent llama3-8b model achieves the best overall results in both settings, with macro-F1 scores of 85.6 in the binary setting and 69.8 in the multi-label setting, notably achieved without any hyper-parameter tuning, unlike the other models. Overall, the fine-tuned models tend to struggle the most with the least represented classes, such as Drugs, Self-harm, and Spam. Interestingly, despite comprising 17.3% of the total entries, the models under perform in the Conflictual category, while they exhibit better performance in the less prevalent Sexual Explicit Content class, which accounts for only 9.7% of tweets. This disparity may indicate the models’ difficulties in identifying subtler features within the Conflictual category.
@@ 8 Limitations
Finally, while we conduct an in-depth analysis using the results of six different models, there is significant room for improvement in terms of analysis and model development. This includes, investigating the performance of models of different architectures and optimising the prompts used [^9] The prompts used in our experiments can be found in Appendix C.3.
@@ 9 Ethics Statement
We also ensure fair treatment of the annotators who labelled the dataset by: 1) compensating them fairly at an average rate of 12\$ per hour, and 2) not sharing or storing personal identification information. As annotator demographics play an important role in the perception of toxicity, following Prabhakaran et al. (2021), we release the data [^10] https://huggingface.co/datasets/cardiffnlp/x_sensitive/blob/main/all_annotations.json, disaggregated by individual annotator labels, while making sure that the demographic information is course enough to prevent deanonymization of the crowd-workers.
@@ Appendix A Data Collection
Table 4 displays the total number of keywords and seed words used in our data collection.
@@ B.2 Annotator Agreement
Table 5: Krippendorff’s alpha within each gender group of coders.
@@
Table 6: Krippendorff’s alpha within each age group of coders.
@@
Age conflictual profanity sex drugs self-harm spam AVG Coders
@@
Gender conflictual profanity sex drugs self-harm spam AVG Coders
@@ C.2 Taxonomies
’self_harm’: ’Self-Harm’,
’self_harm_instructions’: ’Self-Harm’,
’self_harm_intent’: ’Self-Harm’,
@@
’self-harm’: ’Self-Harm’,
@@
’self-harm/intent’: ’Self-Harm’,
’self-harm/instructions’: ’Self-Harm’,
@@
C.3 Prompts
Below are the prompt templates used in the zero- and few-shot settings.
Multi-label
Classify the following sentence in one or more of the categories listed below: 1. self-harm, 2. drugs, 3. conflictual language, 4. spam, 5. profane or bad words, 6 sexually explicit. Answer with the following format: ’ANSWER= 2, 3’ if the sentence contains sensitive content or ’ANSWER= 0’ if the sentence does not contain sensitive content.
Sentence: "{{ message }}"
Binary
Below there is a sentence. I would like to know if it contains sensitive content. We understand the sentence contains sensitive content if it treats topics like self-harm, drugs, if it contains conflictual language, spam, expresses profanity or it is sexually explicit. Does this tweet contain any kind of sensitive content? Please answer with the following format: ’ANSWER= 1’ if the tweet contains sensitive content or ’ANSWER= 0’ if the sentence does not contain sensitive content.
Sentence: "{{ message }}"
C.4 Extended Results